OpenStudy (anonymous):

Convert (x-a)^2+y^2=k into polar co-ordinates

OpenStudy (anonymous):

From 0,0

OpenStudy (anonymous):



OpenStudy (anonymous):

turn it into this form \(r^2=2ar \cos \theta+k-a^2\) ? or .... actually i dont know what do u mean by from 0,0

OpenStudy (anonymous):

 Here is (x-a)^2+y^2=k I'd like the 'centre of polar rotation' (i.e. the dot in the 1st picture) to be at the cartesian co-ordinates of 0,0

OpenStudy (anonymous):

How do you get it into the form that you stated (sorry, I know next to nothing about polar co-ordinate making)

OpenStudy (anonymous):

Converting between polar and Cartesian coordinates \(x^2+y^2=r^2\) \(x=r \cos \theta\) \(y=r\sin \theta\)

OpenStudy (turingtest):

the way you have it\[(x-a)^2+y^2=k\]the radius is\[r=\sqrt k\]so\[x=\sqrt k\cos\theta\]\[y=\sqrt k\sin\theta\]plugging this in gives\[(\sqrt k\cos\theta-a)^2+(\sqrt k\sin\theta)^2=k\]which you can simplify a bit

OpenStudy (turingtest):

I mean\[(\sqrt k\cos\theta-a)^2+(\sqrt k\sin\theta)^2=r^2\]

OpenStudy (turingtest):

I messed this up explanation before, let me try again conversion to polar coordinates is\[x=r\cos\theta\]\[y=r\sin\theta\]and a circle is given by the formula\[x^2+y^2=r^2(\cos^2\theta+\sin^2\theta)=r^2=Constant\]now to shift this circle around the have a center \((h,k)\) all we need to do is make the appropriate adjustment for our changes for x and y:\[x=h+r\cos\theta\]\[y=k+r\sin\theta\]so the circle at center \((h,k)\) in polar is\[(h+r\cos\theta)^2+(k+r\sin\theta)^2=Constant\]so the constant is the square of the actual radius of the circle, and \(r\) is the radius from the origin that traces out the circle as \(\theta\) changes I hope that made sense

OpenStudy (anonymous):

How would you put that into the form r=f(θ)?

OpenStudy (turingtest):

solve for r...

OpenStudy (anonymous):

Rearrange it?

OpenStudy (turingtest):

\[h^2+2rh\cos\theta+r^2+k^2+2rk\sin\theta=C\]I never claimed it to be easy... and with r^2 it may not technically be a function upon solving for r

OpenStudy (turingtest):

this sucker is quadratic in r and you could use the quadratic formula if you wished

OpenStudy (turingtest):

I think the message here is that it kind of defeats the purpose of using polar coordinates to simplify things unless they have a polar symmetry about the origin of some kind even though the graph has a rotational symmetry, the benefit is broken by shifting it and breaking the symmetry of the whole graph resulting in something fairly ugly

OpenStudy (turingtest):

no on the other hand if you leave it parametric and keep the coordinates separated it is fine to use\[x=h+r\cos\theta\]\[y=k+r\sin\theta\]\[\vec f(\theta)=\langle h+r\cos\theta,k+r\sin\theta\rangle\]

OpenStudy (turingtest):

now on the....*

OpenStudy (turingtest):

that is a parametric vector function that traces out the circle

OpenStudy (anonymous):

My problem is that I'm trying to decompose r(θ) into its x and y components to work out the force on an object- damn, you had actually answered that earlier, but I was overcomplicating- Thanks and apologies for dragging this out!

OpenStudy (turingtest):

no problem, a little context helps sometimes :)

OpenStudy (anonymous):

Especially as I am useless at polar equations...

OpenStudy (turingtest):

I am pretty lazy about them too I consult things like this from time to time http://tutorial.math.lamar.edu/Classes/CalcII/PolarCoordinates.aspx